You are a consultant who is advising a monopoly on the optimal pricing strategy. Your analysis has yielded the following information.

  • The marginal cost (MC) is $3.
  • The demand equation is P = 90 - 3Q
  • The total cost (TC)is given by 35 + 3Q
  • The marginal revenue (MR) is given by 90 - 6Q

Based on this information, answer the following questions.  Show FULL calculations!

(a) Following the concepts of profit maximization, what is the profit maximizing quantity for this monopoly?

(b)  Following the concepts of profit maximization, what is the profit maximizing price for this monopoly?

(c)  Following the concepts of profit maximization, what is the monopoly's profit at the profit maximization point?

(d) If the firm produces 10 units, what is the level of fixed costs?

(e) If the firm produces 20 units, what is the level of fixed costs?

(f) If the firm produces 15 units, what is the level of total costs?
Answer #1

Answer:

A. Profit maximizing quantity

MR = MC is condition

90 - 6Q = 3

Q = 87 / 6

Q = 14.5 is profit maximizing quantity

B. Price

Demand curve = 90 - 3Q , putting value Q = 14.5

= 90 - 3(14.5) = 90 - 43.5 = 46.5

P = 46.5 is profit maximizing price

C. Profit

Profit = TR - TC

TR = P*Q = 46.5 * 14.5 = 674.25

TC = 35 + 3Q = 35 + 3(14.5) = 78.5

Profit = 674.25 - 78.5 = 595.75

D Fixed cost for 10 units

- Fixed cost is 35 , AFC will be 3.5

E Fixed cost for 20 units

- Fixed cost is 35, AFC will be 1.75

F Total cost if units are 15

TC = 35 + 3Q = 35 + 3(15) = 80